how do you solve m over 3 = m+4 over 7 for m?
Answer by stanbon(75887)

solve m over 3 = m+4 over 7
------------
Equation:
m/3 = (m+4)/7
Cross-mutiply to get:
7m = 3(m+4)
7m = 3m + 12
4m = 12
m = 3
-------------
Cheers,
Stan H.
================